CH/DOC/94/11

Physics Meeting

Announcement and Agenda

Author: B.R. Fulton

Date: 5 May 1994

Charissa "Physics" Meeting

A full day meeting will be held at Daresbury on Friday
27 May.   The purpose of this meeting is to review the
present state of the analysis of the various experiments
and discuss what needs to be done to complete each.  We
will also hear of the present state of the theoretical
developments which are going on.  It is hoped the meeting
will serve to bring everyone up to date with the results
of the physics programme and also help move forward some of
the analysis and interpretation on the more recent experiments.
Later in the year it is hoped we can have another meeting
at which we will talk about the next phase of experiments
and the directions the theoretical work will develop.
A proposed timetable is attached below.   It includes coverage
of all the outstanding NSF experiments and those from
the first period of running at ANU.  Each speaker should
aim to talk for 10 minutes, allowing 10 minutes for discussion.
The talks in the morning will tend to be reports on
almost completed work where the discussion may focus on the
physics interpretation, while in the afternoon they will be more
of a status report on the analysis where the discussion may
concentrate more on what other things might be looked for
in the data.

0930 Introduction BRF
0940 Texas Data BRF
1000 Higher Excitation region BRF
1020 23Na data DLW
1040 LBL 16O+8Be JP
1100 Coffee
1120 16O target CDJ
1140 High Resolution measurement NC
1200 Band termination SPF
1220 Spin at low energy CDJ
1240 Angular Distn SPF
1300 Lunch
1400 12C on 24Mg SMS
1420 40Ca SMS
1440 28Si/32S NC
1500 7 alpha PS
1520 8Be+8Be MF
1540 O3+ SPC
1600 Status of Theory - where next WDMR
1630 Summary and discussion